OpenStudy (anonymous):

MEDAL Frankie can build a speaker box in 120 minutes. Tisha can build a speaker box 40 minutes faster. If they work together, how many minutes does it take them to build the speaker box?

OpenStudy (anonymous):

answer choices 240 80 48 40

OpenStudy (anonymous):

frankie can do it in 120 minutes, tisha can do it in 80 minutes, together the can do it in \(t\) minutes \[\frac{ 1 }{ 120 }+\frac{ 1 }{ 80 }=\frac{ 1 }{ t }\] multiply by the lcd of 240t \[2t+3t=240\]

OpenStudy (anonymous):

80

OpenStudy (anonymous):

48 sorry